A line has the equation 2x= -6y +5. What is its slope?
Answer by sejackson91(8)

the slope intercept form is y = mx+b where m is the slope and b is the y-intercept
what you need to find is m so rearrange the given formula to y = mx+b
2x = -6y + 5
6y = -2x + 5
y = -(1/3)x + 5/6
so m = -1/3 and that is your slope